OpenStudy (kayders1997):

X is approaching infinity when the limit is equal to xsin(pi/2)

myininaya (myininaya):

you mean evaluate the limit of x sin(pi/2) as x approaches infinity just saying because I don't believe you mean the limit is x sin(pi/2)

myininaya (myininaya):

\[\lim_{x \rightarrow \infty} x \sin(\frac{\pi}{2})\] is this really what you want to evaluate?

myininaya (myininaya):

because this is constant times a large large number

OpenStudy (kayders1997):

Yes it's what I want, it's my homework

OpenStudy (hasankoush):

yeah that should be the format ofthe equation

myininaya (myininaya):

you do know sin(pi/2) is 1 so you have \[\lim_{x \rightarrow \infty} x\]

myininaya (myininaya):

you already know the answer if x approaches infinity, then x approaches ? :)

OpenStudy (hasankoush):

i dont think i should post the answer, right?

OpenStudy (kayders1997):

Omg it's infinity...for some reason I was thinking it was 0 times inifinty

myininaya (myininaya):

and there is really no x inside the sine thing right?

myininaya (myininaya):

and yes you are right @kayders1997

OpenStudy (kayders1997):

Thank you

OpenStudy (kayders1997):

I must have been thinking cos(pi/2)

myininaya (myininaya):

cos(pi/2) is 0 and then your limit would be 0

OpenStudy (kayders1997):

No

myininaya (myininaya):

if it was cos(pi/2) *x as x approaches infinity

OpenStudy (kayders1997):

Maybe idk

myininaya (myininaya):

maybe?

myininaya (myininaya):

what are you referring to with the maybe

OpenStudy (kayders1997):

My teacher said that if it was 0 and inifinty it could be 0 infinity or in between

myininaya (myininaya):

\[\lim_{x \rightarrow \infty} c x=c \cdot \lim_{x \rightarrow \infty}x\]

myininaya (myininaya):

where c is a constant

myininaya (myininaya):

0 is a constant you can pull it outside the limit

OpenStudy (kayders1997):

Okay

myininaya (myininaya):

however if you had this \[\lim_{x \rightarrow \infty} \frac{1}{x} \cdot x\] you cannot do the following \[\lim_{x \rightarrow \infty} \frac{1}{x} \cdot \lim_{x \rightarrow \infty} x\] because one of the limits do not exist (that second one there) and also this form is indeterminate 0*infty

myininaya (myininaya):

but we know x/x=1

myininaya (myininaya):

the so the limit here is 1

myininaya (myininaya):

for this example

myininaya (myininaya):

but you had a constant multiple which could be brought outside the limit always

OpenStudy (kayders1997):

Okay thank you